OpenStudy (anonymous):

If each foot of water in a lake screens out 60% of the light above, what percent of the light passes through 5 feet of water?

OpenStudy (lgbasallote):

\[0.6 \times 0.6 \times 0.6 \times 0.6 \times 0.6\]

OpenStudy (lgbasallote):

because you take the 60% of the one before...you get the logic right?